# Holistic White-light Polyp Classification via Alignment-free Dense Distillation of Auxiliary Optical Chromoendoscopy

### v1

@@ 1 Introduction
Colonoscopy is the gold standard for colorectal cancer (CRC) screening and diagnosis, where accurately distinguishing between adenomatous and hyperplastic polyps is essential [1]. Adenomatous polyps require removal due to their malignant potential, while hyperplastic polyps typically only necessitate periodic monitoring. However, variations in clinical experience can lead to inconsistencies in diagnosis. To enhance diagnostic objectivity, automated polyp classification algorithms for colonoscopy have been extensively studied [2].
@@
To address the accuracy gap and enhance the diagnostic precision of WLI, existing methods [5, 6] have sought to leverage domain alignment strategies, knowledge from NBI to WLI to improve its discriminative capabilities. However, these approaches fundamentally rely on cropped lesion patches for polyp classification, thereby necessitating additional detection steps that may introduce systematic errors and hinder generalizability. Furthermore, by focusing solely on cropped lesion regions and employing global alignment techniques, these methods frequently overlook valuable contextual information (e.g., growth position and color contrast with surrounding tissue) and fail to adequately capture subtle diagnostic vascular information in the NBI domain.
@@ 2.3 Training details
where $\mathcal{L}_{cls}=\mathrm{CE}(P^{w},Y)$ and $\mathcal{L}_{logit}=\|P^{w}-P^{n}\|$. Here, $\text{CE}(\cdot,\cdot)$ denotes the cross-entropy loss, and $P^{w}$ and $P^{n}$ represent the predicted logits of the WLI and NBI classifiers, respectively. The implementation is by PyTorch [10] and the training is on a RTX 4090 GPU with 24GB memory. Weights of ResNet50 [11] pre-trained on ImageNet [12] are loaded as initialization. The input size is set to $448\times 448$, and the batch size is set to $16$. The optimizer is set to Adam [13] with $81e-$ 8 weight decay. The number of epochs is set to $200$ and the initial learning rate is set to $41e-$ 4. We set $T$, $\tau_{1}$, and $\tau_{2}$ as $10$, $0.3$, and $0.7$, respectively, according to the parameter search experiment.
@@ 3.1 Datasets and Evaluation Metrics
We evaluate our method on two datasets: the CPC-Paired dataset [5] and our in-house dataset, both containing misaligned WLI-NBI polyp image pairs. The in-house dataset, collected from a local hospital, was annotated by two experienced colonoscopists. Both datasets consist of two polyp categories: hyperplastic and adenomatous. The CPC-Paired dataset includes $116$ hyperplastic and $307$ adenomatous pairs, while our dataset contains $514$ hyperplastic and $598$ adenomatous pairs.

### v3

@@ 1 Introduction
Colonoscopy is the gold standard for colorectal cancer (CRC) screening and diagnosis, where accurately distinguishing between adenomatous and hyperplastic polyps is essential [1]. Adenomatous polyps require removal due to their malignant potential, while hyperplastic polyps typically only necessitate periodic monitoring. However, variations in clinical experience can lead to inconsistencies in diagnosis. To enhance diagnostic objectivity, automated polyp classification algorithms for colonoscopy have been extensively studied [2, 3, 4].
@@
To address the accuracy gap and enhance the diagnostic precision of WLI, existing methods [7, 8] have sought to leverage domain alignment strategies, knowledge from NBI to WLI to improve its discriminative capabilities. However, these approaches fundamentally rely on cropped lesion patches for polyp classification, thereby necessitating additional detection steps [9, 10] that may introduce systematic errors and hinder generalizability. Furthermore, by focusing solely on cropped lesion regions and employing global alignment techniques, these methods frequently overlook valuable contextual information (e.g., growth position and color contrast with surrounding tissue) and fail to adequately capture subtle diagnostic vascular information in the NBI domain.
@@ 2.3 Training details
where $\mathcal{L}_{cls}=\mathrm{CE}(P^{w},Y)$ and $\mathcal{L}_{logit}=\|P^{w}-P^{n}\|$. Here, $\text{CE}(\cdot,\cdot)$ denotes the cross-entropy loss, and $P^{w}$ and $P^{n}$ represent the predicted logits of the WLI and NBI classifiers, respectively. The implementation is by PyTorch [14] and the training is on a RTX 4090 GPU with 24GB memory. Weights of ResNet-50 [15] pre-trained on ImageNet [16] are loaded as initialization. The input size is set to $448\times 448$, and the batch size is set to $16$. The optimizer is set to Adam [17] with $81e-$ 8 weight decay. The number of epochs is set to $200$ and the initial learning rate is set to $41e-$ 4. We set $T$, $\tau_{1}$, and $\tau_{2}$ as $10$, $0.3$, and $0.7$, respectively, according to the parameter search experiment.
@@ 3.1 Datasets and Evaluation Metrics
We evaluate our method on two datasets: the CPC-Paired dataset [7] and our in-house dataset, both containing misaligned WLI-NBI polyp image pairs. The in-house dataset, collected from a local hospital, was annotated by two experienced colonoscopists. Both datasets consist of two polyp categories: hyperplastic and adenomatous. The CPC-Paired dataset includes $63$ hyperplastic and $102$ adenomatous pairs, while our in-house dataset contains $514$ hyperplastic and $598$ adenomatous pairs.
@@ 4 Conclusion
{credits}
4.0.1 Acknowledgements
This work was supported in part by National Natural Science Foundation of China (Grant No. 62202189), Key R&D Program of Hubei Province of China (No. 2023BCB003), and Changzhou United Imaging Surgical Co., Ltd. Thanks to Professor Mei Liu’s team for providing the in-house dataset used in this work at the Department of Gastroenterology, Tongji Medical College, Huazhong University of Science and Technology.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.